# Single-Shot Readout and Weak Measurement of a Tin-Vacancy Qubit in Diamond

### v1

@@ Abstract
The negatively charged tin-vacancy center in diamond (${}^{-}$) is an emerging platform for building the next generation of long-distance quantum networks. This is due to the ${}^{-}$’s favorable optical and spin properties including bright emission, insensitivity to electronic noise, and long spin coherence times at temperatures above 1 Kelvin. Here, we demonstrate measurement of a single ${}^{-}$ electronic spin with a single-shot readout fidelity of 87.4%, which can be further improved to 98.5% by conditioning on multiple readouts. We show this performance is compatible with rapid microwave spin control, demonstrating that the trade-off between optical readout and spin control inherent to group-IV centers in diamond can be overcome for the ${}^{-}$. Finally, we use weak quantum measurement to study measurement induced dephasing; this illuminates the fundamental interplay between measurement and decoherence in quantum mechanics, and makes use of the qubit’s spin coherence as a metrological tool. Taken together, these results overcome an important hurdle in the development of the ${}^{-}$ based quantum technologies, and in the process, develop techniques and understanding broadly applicable to the study of solid-state quantum emitters.
@@ I Introduction
Color center qubits have been recognized as an advantageous platform for the realization of quantum technologies, and in particular quantum networks [1], due to their efficient spin-photon interface, long spin coherence times, and compatibility with nanophotonics [2, 3, 4]. An outstanding challenge is to scale quantum networks to include more nodes, greater distance between nodes, and further complexity including error corrected registers of qubits within each node [5].
@@
The path to solving these many challenges is specific to the choice of quantum platform and, in this case, choice of color center. Today, state-of-the art quantum networks consist of three nodes, where each node is based on a single nitrogen-vacancy center (${}^{-}$) in diamond [6, 7]. However, the ${}^{-}$ is not the optimal qubit for future quantum networks because of its sensitivity to electrical noise, and a low probability of emission into its zero-phonon line (ZPL), which reduces entanglement generation rate.
@@
Of countless optically active solid-state atomic defects including rare earth ions [8, 9] and defects in silicon carbide [10], group-IV centers in diamond have emerged as promising candidates to advance quantum networks. These qubits have a centrosymmetric structure, which renders a first-order insensitivity to electrical noise, allowing for relative optical stability even within nanophotonic structures [2]. Advantageously, these color centers also have high quantum efficiency and strong emission into their ZPL, promising high rates of entanglement generation. Among them, the silicon-vacancy center (${}^{-}$) [11] is the most developed, with demonstrations of long spin coherence times and single-shot electron spin readout [12], integration with nanophotonic cavities [13, 14, 15, 16, 17], and demonstration of a two-node quantum network [18].
@@
However, due to its spin-orbit induced ground state splitting of 50 GHz, the ${}^{-}$ requires operation at millikelvin temperatures to avoid thermal decoherence [11]. Even then, state-of-the-art demonstrations use highly strained ${}^{-}$’s to increase this splitting and reduce drive induced heating [15, 16, 17]. To avoid temperature related issues, the tin-vacancy center (${}^{-}$) in diamond has emerged as a favorable alternative due to its larger minimum ground state splitting of $\approx 820\,\mathrm{GHz}$ [19, 20, 21, 22, 23, 24]. This allows for coherent spin control at several Kelvin [25, 26, 27], where exponentially more cooling power is available.
@@
Recent advances of the ${}^{-}$ platform include incorporation with nanophotonics [28, 29, 30, 31, 32, 33, 34], high-fidelity generation of single photons [31, 35], nuclear spin control and single-shot nuclear spin readout [32] enabled by a large hyperfine coupling [36], and high-fidelity microwave spin control using both moderately strained [26] and highly strained [27, 24] centers.
@@
However, recent work on ${}^{-}$ spin control has illuminated a complicated relationship between the performance of spin polarization, coherent microwave spin control, and spin readout as functions of strain and the orientation of the applied magnetic field [26, 27, 37, 38, 24]. In particular, there is a general trade-off between optimizing for high-fidelity microwave spin control using high strain and certain magnetic field orientations, and optimizing readout using low strain and a magnetic field aligned with the spin dipole axis. To-date, the only published demonstration of single-shot readout of an ${}^{-}$’s electronic spin uses an unstrained center not favorable for coherent spin control [39], and the only demonstrations of coherent spin control use readout far from the single-shot regime [25, 26, 27, 24]. This brings into question if rapid spin control and high-fidelity readout are compatible for the ${}^{-}$, as has previously been shown for the ${}^{-}$ [12].
@@
In this article, by precise study of readout performance and associated trade-offs (Fig. 1), we answer this question favorably and further our understanding about both the ${}^{-}$ platform and, in general, about the quantum measurement of solid-state emitters. We demonstrate measurement of a single ${}^{-}$ electronic spin with a single-shot readout fidelity of $87.4\%$ at the same operating conditions as rapid microwave spin control. We also report a conditional readout fidelity of $98.5\%$ by conditioning on the outcome of two consecutive measurements. We achieve this performance using an overall measurement efficiency of $\approx 0.1\%$, implying that near-unit fidelity is achievable in future nanophotonic devices that have much greater efficiency. Finally, we use a combination of coherent spin control and weak quantum measurement to study measurement induced dephasing — both affirming the basic science of quantum measurement, and, using the qubit’s spin superposition as a resource to benchmark its interaction with light and characterize the measurement apparatus. The understanding and methods developed in this work will advance ${}^{-}$ based quantum technologies, will enable the use of an ${}^{-}$’s long spin coherence time as a resource for quantum memory, and will have broad application to the study and measurement of a wide class of solid-state quantum systems.
@@
Figure 1: Schematic. A tin-vacancy (${}^{-}$) qubit consists of a tin atom replacing two carbon atoms in a diamond lattice. Here, we manipulate the spin of this atom-like system under application of a static magnetic field, and using a combination of optical and microwave control pulses. This leads to spin-dependent photoluminescence, which we measure using a single photon counter. The presence or absence of “clicks” on the detector is used to determine the qubit’s spin state with high fidelity. Even so, some emission is lost via different channels before detection; this loss is parameterized by a beamsplitter with transmittivity $\eta$ placed between the qubit and an idealized detector. Here $\eta$ is the overall measurement efficiency.
@@
Figure 2: Spin-selective state initialization and optical and microwave control. (a) Simplified energy level diagram of the tin-vacancy qubit (${}^{-}$) in diamond. Electron spin ground states $\ket{\downarrow}$ and $\ket{\uparrow}$ are separated by the qubit frequency $\omega_{q}/2\pi$. Red and orange lines illustrate optical transitions to the excited states |$\ket{A}$ and |$\ket{B}$. These transitions are either spin-preserving (A1 and B2), or spin-flipping (A2 and B1). (b) The spin, with magnetic dipole $\vec{\mu}$, interacts with a magnetic field $\vec{B}$ via the Zeeman effect. Here, $\vec{B}$ is oriented at the angle $\zeta$ in the plane defined by the $\hat{x}$ and $\hat{z}$ coordinates. The spin is oriented at azimuthal angle $\varphi$ from this plane. (c) Photoluminescence excitation (PLE) measurement of the A1 and B2 transitions vs $\zeta$, fixing $|\vec{B}|=180\,\mathrm{mT}$. Detuning is measured from 619.13972 nm (484.20808 THz). (d) Driving a spin-preserving transition polarizes the spin at rate $\kappa$ (data shown for driving A1). This rate is also dependent on $\zeta$. (e) Cyclicity (the ratio of the probabilities of spin-preserving to spin-flipping decay) vs. $\zeta$, determined from a fit of $\kappa$ applied to Eq. 1 in the high power limit. Here, cyclicity changes by three orders-of-magnitude depending on $\zeta$, up to $\chi\approx 2.5\times 10^{3}$. Cyclicity is expected to diverge when $\vec{B}$ and $\vec{\mu}$ align; here it is limited by the minimum misalignment $\varphi$, set to $\varphi=10\degree$ for a model which follows our data (solid black line). (c) and (e) share an x-axis. (f) Rabi oscillations between the ground spin states under microwave driving, demonstrating an 80 ns $\pi$-pulse time at $\omega_{q}/2\pi=3.677\,\mathrm{GHz}$. (g) Microwave Rabi rate vs. $\zeta$. At each angle, the amplitude $|\vec{B}|$ is tuned so that $\omega_{q}/2\pi=4.00\,\mathrm{GHz}$. Data are compared to a model [26] that uses a microwave field of amplitude 0.6 mT oriented perpendicular to $\vec{\mu}$.
@@ II.1 Optical transitions
Of crucial importance to the measurement of any qubit is its Hamiltonian. For group-IV color centers in diamond, the Hamiltonian [40] is set by both strain of the diamond lattice and by $\vec{B}$, the external magnetic field. Here we study the same ${}^{-}$ used in Ref. [26], which has a ground state splitting of 903 GHz, larger than that of an unstrained ${}^{-}$. This is in a regime of “moderate” strain, where the ground-state strain ($\approx 355\,\mathrm{GHz}$) is non-negligible but less than the spin orbit coupling.
To understand the importance of magnetic field orientation, we first consider the ${}^{-}$’s two lowest energy levels $\ket{\downarrow}$ and $\ket{\uparrow}$ in the ground state manifold, and the two lowest energy levels |$\ket{A}$ and |$\ket{B}$ in the excited state manifold, Fig. 2a. Both sets of levels split with magnetic field due to the Zeeman effect, which can lead to four distinct transitions as shown in Fig. 2a: A1 and B2 (“spin-preserving”), and A2 and B1 (“spin-flipping”) [11]. The detuning between these transitions is highly dependent on the alignment between the magnetic field $\vec{B}$ and the magnetic spin dipole $\vec{\mu}$, parameterized by the angle $\zeta$, which is the angle of the vector $\vec{B}$ in the plane defined by our two-axis magnet with coils along the $\hat{x}$ and $\hat{z}$ directions (see Fig. 2b). The dipole axis is defined by the crystallographic symmetry axis that connects two missing carbon atoms and an interstitial tin vacancy (see Fig. 1). The angle $\varphi$ is a measure of misalignment between the dipole and magnetic field plane, and we have tried to minimize this angle by physical rotation of the sample.
The A1 and B2 transitions are characterized using photoluminescence excitation (PLE) spectroscopy, Fig. 2c, by sweeping $\zeta$ for a magnetic field of fixed amplitude $|\vec{B}|=180\,\mathrm{mT}$. Data in Fig. 2 are modelled by the Hamiltonian described in Table. I of Ref. [24]. See Appendix B.1 for details, including discussion of free parameters and modifications to the ${}^{-}$ Hamiltonian reported by our previous work, Ref. [26].
@@ II.2 Cyclicity
In addition to setting the detuning between the A1 and B2 transitions, the magnetic field orientation controls the cyclicity; a crucial parameter for understanding spin measurement. When pumped to |$\ket{A}$ or |$\ket{B}$, the ${}^{-}$ subsequently relaxes to its ground state in a manner which can preserve the spin, i.e. by decay via the A1 or B2 transitions with probabilities $P_{\mathrm{A1}}$ or $P_{\mathrm{B2}}$, respectively. Alternatively, its spin can flip: i.e. by decay via the A2 or B1 transitions with probabilities $P_{\mathrm{A2}}$ and $P_{\mathrm{B1}}$, respectively. Cyclicity $\chi=P_{\mathrm{A1}}/P_{\mathrm{A2}}=P_{\mathrm{B2}}/P_{\mathrm{B1}}$ is the ratio of the probability of spin-preserving decay to spin-flipping decay.
@@
Cyclicity determines how many photons are emitted, corresponding to the maximum allowed number that can be collected during readout, and the optimal readout duration. For a given excitation power, a higher $\chi$ indicates more photons can be collected before the spin polarizes and goes “dark” (no longer emits photons). However, in trade-off, a longer duration of optical drive is required to polarize the spin state. A smaller $\chi$ will lead to faster polarization, but will limit the number of emitted photons and the useful duration of acquisition.
@@
For example, if the ${}^{-}$ is prepared in the $\ket{\downarrow}$ state given by density matrix $\hat{\rho}=\ket{\downarrow}\bra{\downarrow}$ | and driven on A1 starting at time $t=0$, it will polarize into $\ket{\uparrow}$ at rate $\kappa$ such that $\hat{\rho}(t)=e^{-\kappa t}\ket{\downarrow}\bra{\downarrow}+(1-e^{-\kappa t})\ket{\uparrow}\bra{\uparrow}$ |. Here,
$\kappa=\frac{R}{\chi+1},$ (1)
is the polarization rate and,
@@
captures an effective optical pumping rate to |$\ket{A}$, which depends on drive power $p$, on-resonance saturation power $p_{\mathrm{sat}}$, and detuning $\delta$ of the drive from the transition. Eq. 1 follows from the optical Bloch equations [41], see Appendix A.1 for details. For the ${}^{-}$, $\gamma\approx(4.5\pm 0.2\,\mathrm{ns})^{-1}=2\pi\times(35\pm 1.6\,\mathrm{MHz})$ is the optical decay rate [22].
@@
Intuitively, at $p/p_{\mathrm{sat}}\gg 1$ and $\delta\lesssim\gamma$, the rate of excitation to the higher state is much faster than relaxation, thus $\kappa\approx\gamma/2\chi$ is independent of power but is proportional to the saturated optical decay rate and inversely proportional to cyclicity. At $p/p_{\mathrm{sat}}\ll 1$, however, the excitation rate is much slower than relaxation, and $\kappa\approx(\gamma/2\chi)\times(p/p_{\mathrm{sat}})$ is also proportional to power.
@@
We experimentally measure $\kappa$ by binning the time-tagged photon counts collected during resonant excitation. The number of collected counts decays exponentially: $(\Gamma_{\mathrm{max}}-\Gamma_{\mathrm{min}})e^{-\kappa t}+\Gamma_{\mathrm{min}}$, and data in Fig. 2d is fit to this exponential model to determine $\kappa$. Here $t=0$ is the onset of excitation, $\Gamma_{\mathrm{max}}$ is the maximum count rate, and $\Gamma_{\mathrm{min}}$ is the mean count rate at $t\gg 0$. The term $(\Gamma_{\mathrm{max}}-\Gamma_{\mathrm{min}})e^{-\kappa t}$ captures signal from ${}^{-}$ emission, and the background $\Gamma_{\mathrm{min}}$ arises from noise dominated by unwanted scatter of excitation light into collection and detector dark counts. Cyclicity $\chi\approx\gamma/2\kappa$ is determined from $\kappa$ via Eq. 1. For data in Fig. 2 we operate at negligible detuning (the drive frequency is re-calibrated at each point) and at powers far above saturation.
@@
Both $\kappa$ and $\chi$ are highly dependent on the alignment between $\vec{B}$ and $\vec{\mu}$. This is characterized in Fig. 2e by measuring $\chi$ vs. the angular orientation of $\vec{B}$ along angle $\zeta$. Cyclicity changes by three orders of magnitude during this sweep. At $\zeta=147\degree$ near where alignment is maximum, we measure $\kappa=(20.4\pm 0.30\,\mathrm{\mu s})^{-1}=49.02\pm 0.72\,\mathrm{kHz}$ corresponding to $\chi=2244\pm 108$. At similar power and at $\zeta=53\degree$ where alignment is near minimum, we measure $\kappa=(78.5\pm 0.002\,\mathrm{ns})^{-1}=12.74\pm 0.0004\,\mathrm{MHz}$ and determine $\chi=8.6\pm 0.4$. The difference between $\zeta$ values corresponding to maximum and minimum cyclicity differs from exactly $90\degree$; we postulate this is due to slight miscalibration of one or both magnetic coils, such that amplitude $|\vec{B}|$ is also changing slightly with $\zeta$.
@@
When $\vec{B}$ and $\vec{\mu}$ are perfectly aligned, spin-flipping relaxation becomes entirely forbidden, and cyclicity is expected to diverge ($\varphi=0$ model in Fig. 2e). The extent to which it does not indicates remaining misalignment, e.g. the spin dipole $\vec{\mu}$ is rotated by azimuthal angle $\varphi$ relative to the plane defined by the axes $\hat{x}$ and $\hat{z}$, Fig. 2b. Our data follow a model using a misalignment of $\varphi=10\degree$ (solid black line, Fig. 2e). This suggests cyclicity can be increased by fine tuning sample rotation and/or by using a three-axis vector magnet.
@@
Cyclicity is minimal when $\vec{B}\cdot\vec{\mu}\approx 0$. However, the models in Fig. 2e (based on the model in Ref. [26]) show a slight increase in cyclicity around maximal misalignment. This feature is due to the re-definition of the spin’s quantization axis along the direction of strain. Finally, we notice that for some of the angles the data differs from the model where points of lowest cyclicity occur near the angles where the A1 and B2 transitions become degenerate. We postulate this effect could be due to a coupling interaction between these transitions indicative of a previously unexplored term in the ${}^{-}$ Hamiltonian; further investigation is necessary to understand this phenomenon.

### v2

@@ Abstract
The negatively charged tin-vacancy center in diamond (SnV-) is an emerging platform for building the next generation of long-distance quantum networks. This is due to the SnV-’s favorable optical and spin properties including bright emission, insensitivity to electronic noise, and long spin coherence times at temperatures above 1 Kelvin. Here, we demonstrate measurement of a single SnV- electronic spin with a single-shot readout fidelity of 87.4%, which can be further improved to 98.5% by conditioning on multiple readouts. In the process, we develop understanding of the relationship between strain, magnetic field, spin readout, and microwave spin control. We show that high-fidelity readout is compatible with rapid microwave spin control, demonstrating a favorable parameter regime for use of the SnV- center as a high-quality spin-photon interface. Finally, we use weak quantum measurement to study measurement-induced dephasing; this illuminates the fundamental interplay between measurement and decoherence in quantum mechanics, and provides a universal method to characterize the efficiency of color center spin readout. Taken together, these results overcome an important hurdle in the development of SnV--based quantum technologies, and in the process, develop techniques and understanding broadly applicable to the study of solid-state quantum emitters.
@@ I Introduction
Color center qubits have been recognized as an advantageous platform for the realization of quantum technologies, and, in particular quantum networks [1], due to their efficient spin-photon interface, long spin coherence times, and compatibility with nanophotonics [2, 3, 4]. An outstanding challenge is to scale quantum networks to include more nodes, greater distance between nodes, and further complexity including error corrected registers of qubits within each node [5].
@@
The path to solving these many challenges is specific to the choice of quantum platform and, in this case, choice of color center. Today, state-of-the-art quantum networks consist of three nodes, where each node is based on a single nitrogen-vacancy center (NV-) in diamond [6, 7]. However, the NV- is not the optimal qubit for future quantum networks because of its sensitivity to electrical noise and a low probability of emission into its zero-phonon line (ZPL), which reduces entanglement generation rate.
@@
Of countless optically active solid-state atomic defects including rare earth ions [8, 9] and defects in silicon carbide [10], group-IV centers in diamond have emerged as promising candidates to advance quantum networks. These qubits have a centrosymmetric structure, which renders a first-order insensitivity to electrical noise, allowing for relative optical stability even within nanophotonic structures [2]. Advantageously, these centers also have high quantum efficiency ($80\%$ for the SnV- [11]) and strong emission into their ZPL (Debye-Waller factors of approximately $60\%$ [12]). This strong, coherent emission promises high rates of entanglement generation compared to quantum networks based on the NV-, which has a Debye-Waller factor of only $3\%$ [13]. Among group IV’s, the silicon-vacancy center (SiV-) [14] is the most developed, with demonstrations of long spin coherence times and single-shot electron spin readout [15], integration with nanophotonic cavities [16, 17, 18, 19, 20, 21, 22], and demonstration of a two-node quantum network [23].
@@
However, because the SiV- has the smallest spin-orbit-induced ground state splitting (50 GHz) of all group-IV centers, it is also the most naturally susceptible to decoherence due to thermal excitation out of the spin subspace. State-of-the-art SiV--based experiments, thus, use highly strained SiV-’s to increase this splitting and reduce susceptibility to drive-induced heating, and, furthermore, operate at millikelvin temperatures in dilution refrigerators [19, 20, 21, 22]. To avoid challenges associated with these requirements, the tin-vacancy center (SnV-) in diamond has emerged as a favorable alternative due to its larger minimum ground state splitting of approximately $820\,\mathrm{GHz}$ [24, 11, 25, 26, 27, 28]. This allows for coherent spin control at several degrees Kelvin [29, 30, 31], where exponentially more cooling power is available.
@@
Recent advances of the SnV- platform include incorporation with nanophotonics [32, 33, 34, 35, 36, 37, 38], high-fidelity generation of single photons [35, 39], nuclear spin control and single-shot nuclear spin readout [36] enabled by a large hyperfine coupling [40], and high-fidelity microwave spin control using both moderately strained [30] and highly strained [31, 28] centers.
@@
However, recent work on SnV- spin control has illuminated a complicated relationship between the performance of spin polarization, coherent microwave spin control, and spin readout as functions of strain and the orientation of the applied magnetic field [30, 31, 41, 42, 28]. In particular, there is a general trade-off between optimizing for high-fidelity microwave spin control using high strain and certain magnetic field orientations, and optimizing readout using low strain and a magnetic field aligned with the spin dipole axis. To date, the only published demonstration of single-shot readout of an SnV-’s electronic spin uses an unstrained center not favorable for coherent spin control [43], and the only demonstrations of coherent spin control use readout far from the single-shot regime [29, 30, 31, 28]. This brings into question if rapid spin control and high-fidelity readout are compatible for the SnV-, as has previously been shown for the SiV- [15].
@@
In this article, by precise study of readout performance and associated trade-offs (Fig. 1), we answer this question favorably and further our understanding about both the SnV- platform and, in general, about the quantum measurement of solid-state emitters. We demonstrate measurement of a single SnV- electronic spin with a single-shot readout fidelity of $87.4\%$ at the same operating conditions as rapid microwave spin control. We also report a conditional readout fidelity of $98.5\%$ by conditioning on the outcome of two consecutive measurements. We achieve this performance using an overall measurement efficiency of approximately $0.1\%$, implying that near-unit fidelity is achievable in future nanophotonic devices that have much greater efficiency. Finally, we use a combination of coherent spin control and weak quantum measurement to study measurement-induced dephasing — both affirming the basic science of quantum measurement and using the qubit’s spin superposition as a resource to benchmark its interaction with light and characterize the measurement apparatus. The understanding and methods developed in this work advance SnV- based quantum technologies, enable the use of an SnV-’s long spin coherence time as a resource for quantum memory, and have broad application to the study and measurement of a wide class of solid-state quantum systems.
@@
Figure 1: Schematic. A tin-vacancy (SnV-) qubit consists of a tin atom replacing two carbon atoms in a diamond lattice. Here, we manipulate the spin of this atomlike system under application of a static magnetic field and using a combination of optical and microwave control pulses. This leads to spin-dependent photoluminescence, which we measure using a single-photon counter. The presence or absence of “clicks” on the detector is used to determine the qubit’s spin state with high fidelity. Even so, some emission is lost via different channels before detection; this loss is parameterized by a beam splitter with transmittivity $\eta$ placed between the qubit and an idealized detector. Here, $\eta$ is the overall measurement efficiency.
@@
Figure 2: Spin-selective state initialization and optical and microwave control. (a) Simplified energy level diagram of the tin-vacancy qubit (SnV-) in diamond. Electron spin ground states $\ket{\downarrow}$ and $\ket{\uparrow}$ are separated by the qubit frequency $\omega_{q}/2\pi$. Red and orange lines illustrate optical transitions to the excited states |$\ket{A}$ and |$\ket{B}$, respectively. These transitions are either spin preserving (A1 and B2), or spin flipping (A2 and B1). (b) The spin, with magnetic dipole $\vec{\mu}$, interacts with a magnetic field $\vec{B}$ via the Zeeman effect. Here, $\vec{B}$ is oriented at the angle $\zeta$ in the plane defined by the $\hat{x}$ and $\hat{z}$ coordinates. The spin is oriented at azimuthal angle $\varphi$ from this plane. (c) Photoluminescence excitation (PLE) measurement of the A1 and B2 transitions vs $\zeta$, fixing $|\vec{B}|=180\,\mathrm{mT}$. Detuning is measured from 619.13972 nm (484.20808 THz). (d) Driving a spin-preserving transition polarizes the spin at rate $\Gamma_{p}$ (data shown for driving A1). This rate is also dependent on $\zeta$. (e) Cyclicity (the ratio of the probabilities of spin-preserving to spin-flipping decay) vs $\zeta$, determined from a fit of $\Gamma_{p}$ applied to Eq. 1 in the high-power limit. Here, cyclicity changes by 3 orders of magnitude depending on $\zeta$, up to $\Lambda\approx 2.5\times 10^{3}$. Cyclicity is expected to diverge when $\vec{B}$ and $\vec{\mu}$ align; here it is limited by the minimum misalignment $\varphi$, set to $\varphi=10\degree$ for a model which follows our data (solid black line). (c) and (e) share an x axis. (f) Rabi oscillations between the ground spin states under microwave driving, demonstrating an 80 ns $\pi$-pulse time at $\omega_{q}/2\pi=3.677\,\mathrm{GHz}$. (g) Microwave Rabi frequency vs $\zeta$. At each angle, the amplitude $|\vec{B}|$ is tuned so that $\omega_{q}/2\pi=4.00\,\mathrm{GHz}$. Data are compared to a model [30] that uses a microwave field of amplitude 0.6 mT oriented perpendicular to $\vec{\mu}$.
@@ II.1 Optical transitions
Of crucial importance to the measurement of any qubit is its Hamiltonian. For group-IV color centers in diamond, the Hamiltonian [44] is set by both strain of the diamond lattice and by $\vec{B}$, the external magnetic field. Here, we study the same SnV- used in Ref. [30], which has a ground state splitting of 903 GHz, larger than that of an unstrained SnV-. This is in a regime of “moderate” strain, where the ground state strain (approximately $355\,\mathrm{GHz}$) is non-negligible but less than the spin-orbit coupling.
To understand the importance of magnetic field orientation, we first consider the SnV-’s two lowest energy levels $\ket{\downarrow}$ and $\ket{\uparrow}$ in the ground state manifold and the two lowest energy levels |$\ket{A}$ and |$\ket{B}$ in the excited state manifold, Fig. 2a. Both sets of levels split with magnetic field due to the Zeeman effect, which can lead to four distinct transitions as shown in Fig. 2a: A1 and B2 (“spin-preserving”), and A2 and B1 (“spin-flipping”) [14]. The detuning between these transitions is highly dependent on the alignment between the magnetic field $\vec{B}$ and the magnetic spin dipole $\vec{\mu}$, parameterized by the angle $\zeta$, which is the angle of the vector $\vec{B}$ in the plane defined by our two-axis magnet with coils along the $\hat{x}$ and $\hat{z}$ directions (see Fig. 2b). The dipole axis is defined by the crystallographic symmetry axis that connects two missing carbon atoms and an interstitial tin vacancy (see Fig. 1). The angle $\varphi$ is a measure of misalignment between the dipole and magnetic field plane, and we have tried to minimize this angle by physical rotation of the sample.
The A1 and B2 transitions are characterized using photoluminescence excitation (PLE) spectroscopy, Fig. 2c, by sweeping $\zeta$ for a magnetic field of fixed amplitude $|\vec{B}|=180\,\mathrm{mT}$. Data in Fig. 2 are modeled by the Hamiltonian described in Table. I in Ref. [28]. See Appendix B.1 for details, including discussion of free parameters and modifications to the SnV- Hamiltonian reported by our previous work, Ref. [30].
@@ II.2 Cyclicity
In addition to setting the detuning between the A1 and B2 transitions, the magnetic field orientation controls the cyclicity; a crucial parameter for understanding spin measurement. When pumped to |$\ket{A}$ or |$\ket{B}$, the SnV- subsequently relaxes to its ground state in a manner which can preserve the spin, i.e. by decay via the A1 or B2 transitions with probabilities $P_{\mathrm{A1}}$ or $P_{\mathrm{B2}}$, respectively. Alternatively, its spin can flip: i.e. by decay via the A2 or B1 transitions with probabilities $P_{\mathrm{A2}}$ and $P_{\mathrm{B1}}$, respectively. Cyclicity $\Lambda=P_{\mathrm{A1}}/P_{\mathrm{A2}}=P_{\mathrm{B2}}/P_{\mathrm{B1}}$ is the ratio of the probability of spin-preserving decay to spin-flipping decay.
@@
Cyclicity determines how many photons are emitted, corresponding to the maximum allowed number that can be collected during readout and the optimal readout duration. For a given excitation power, a higher $\Lambda$ indicates more photons can be collected before the spin polarizes and goes “dark” (no longer emits photons). However, in trade-off, a longer duration of optical drive is required to polarize the spin state. A smaller $\Lambda$ will lead to faster polarization, but will limit the number of emitted photons and the useful duration of acquisition.
@@
For example, if the SnV- is prepared in the $\ket{\downarrow}$ state given by density matrix $\hat{\rho}=\ket{\downarrow}\bra{\downarrow}$ | and driven on A1 starting at time $t=0$, it will polarize into $\ket{\uparrow}$ at rate $\Gamma_{p}$ such that $\hat{\rho}(t)=e^{-\Gamma_{p}t}\ket{\downarrow}\bra{\downarrow}+(1-e^{-\Gamma_{p}t})\ket{\uparrow}\bra{\uparrow}$ |. Here,
$\Gamma_{p}=\frac{R}{\Lambda+1}$ (1)
is the polarization rate and
$R=\frac{\gamma}{2}\left(\frac{p/p_{\mathrm{sat}}}{1+p/p_{\mathrm{sat}}+(\frac{\delta}{\gamma/2})^{2}}\right)$ (2)
@@
captures an effective optical pumping rate to |$\ket{A}$, which depends on drive power $p$, on-resonance saturation power $p_{\mathrm{sat}}$, and detuning $\delta$ of the drive from the transition. Equation 1 follows from the optical Bloch equations [45]; see Appendix A.1 for details. For the SnV-, $\gamma\approx(4.5\pm 0.2\,\mathrm{ns})^{-1}=2\pi\times(35\pm 1.6\,\mathrm{MHz})$ is the optical decay rate [26].
@@
Intuitively, at $p/p_{\mathrm{sat}}\gg 1$ and $\delta\lesssim\gamma$, the rate of excitation to the higher state is much faster than relaxation; thus, $\Gamma_{p}\approx\gamma/2\Lambda$ is independent of power but is proportional to the saturated optical decay rate and inversely proportional to cyclicity. At $p/p_{\mathrm{sat}}\ll 1$, however, the excitation rate is much slower than relaxation, and $\Gamma_{p}\approx(\gamma/2\Lambda)\times(p/p_{\mathrm{sat}})$ is also proportional to power.
@@
We experimentally measure $\Gamma_{p}$ by binning the time-tagged photon counts collected during resonant excitation. The number of collected counts decays exponentially: $(a-b)e^{-\Gamma_{p}t}+b$, and data in Fig. 2d are fit to this exponential model to determine $\Gamma_{p}$. Here, $t=0$ is the onset of excitation, $a$ is the maximum count rate, and $b$ is the mean count rate at $t\gg 0$. The term $(a-b)e^{-\Gamma_{p}t}$ captures signal from SnV- emission, and the background $b$ arises from noise dominated by unwanted scatter of excitation light into collection and detector dark counts. Cyclicity $\Lambda\approx\gamma/2\Gamma_{p}$ is determined from $\Gamma_{p}$ via Eq. 1. For data in Fig. 2 we operate at negligible detuning (the drive frequency is recalibrated at each point) and at powers far above saturation.
@@
Both $\Gamma_{p}$ and $\Lambda$ are highly dependent on the alignment between $\vec{B}$ and $\vec{\mu}$. This is characterized in Fig. 2e by measuring $\Lambda$ vs the angular orientation of $\vec{B}$ along angle $\zeta$. Cyclicity changes by 3 orders of magnitude during this sweep. At $\zeta=147\degree$, near where alignment is maximum, we measure $\Gamma_{p}=(20.4\pm 0.30\,\mathrm{\mu s})^{-1}=49.02\pm 0.72\,\mathrm{kHz}$ corresponding to $\Lambda=2244\pm 108$. At similar power and at $\zeta=53\degree$, where alignment is near minimum, we measure $\Gamma_{p}=(78.5\pm 0.002\,\mathrm{ns})^{-1}=12.74\pm 0.0004\,\mathrm{MHz}$ and determine $\Lambda=8.6\pm 0.4$. The difference between $\zeta$ values corresponding to maximum and minimum cyclicity is $94\degree$, differing from the expected value of exactly $90\degree$; we postulate that this is due to slight miscalibration of one or both magnetic coils, such that amplitude $|\vec{B}|$ is also changing slightly with $\zeta$.
@@
When $\vec{B}$ and $\vec{\mu}$ are perfectly aligned, spin-flipping relaxation becomes entirely forbidden, and cyclicity is expected to diverge ($\varphi=0$ model in Fig. 2e). The extent to which it does not indicates remaining misalignment; e.g., the spin dipole $\vec{\mu}$ is rotated by azimuthal angle $\varphi$ relative to the plane defined by the axes $\hat{x}$ and $\hat{z}$, Fig. 2b. Our data follow a model using a misalignment of $\varphi=10\degree$ (solid black line, Fig. 2e). This suggests that cyclicity can be increased by fine-tuning sample rotation and/or by using a three-axis vector magnet.
@@
Cyclicity is minimal when $\vec{B}\cdot\vec{\mu}\approx 0$. However, the models in Fig. 2e (based on the model in Ref. [30]) show a slight increase in cyclicity around maximal misalignment. This feature is due to the redefinition of the spin’s quantization axis along the direction of strain. Finally, we notice that for some of the angles the data differ from the model where points of lowest cyclicity occur near the angles where the A1 and B2 transitions become degenerate. We postulate that this effect could be due to a coupling interaction between these transitions indicative of a previously unexplored term in the SnV- Hamiltonian; further investigation is necessary to understand this phenomenon.